A temperature of 200°F is equivalent to approximately
A. 37.8°C.
B. 93.3°C.
C. 232°C.
D. 840°C

Answer by drk(1908)

we have the equation
F = (9/5)C + 32
and F = 200
so, we get
200 = (9/5)C + 32
subtract 32
168 = (9/5)C
multiply by (5/9) to get
C = 93.333
answer is (B)